OpenStudy (anonymous):

Which of the following shows the right side of the equation after applying a reasonable first step to solve -h/2 + -2/3 = 5/8? 5/8 + 5/8 5/8 + 2/3 5/8 - 5/8 5/8 - 2/3

OpenStudy (dangerousjesse):

B :) And the answer would be \(\large h=\frac{-31}{12}\)

OpenStudy (anonymous):

Thank you! Help with more?

OpenStudy (anonymous):

Human hair grows at a rate of approximately 3/20 cm per year. How many centimeters does it grow in 1/2 of a year? 7/20 centimeters 3/40 centimeters 13/20 centimeters 3/10 centimeters

OpenStudy (dangerousjesse):

\(\huge \color{blue}{\frac{3}{20}\times \frac{1}{2}}\)

OpenStudy (anonymous):

Huh?

OpenStudy (dangerousjesse):

Just multiply :) You'll get \(\large \color{darkblue}{\frac{3}{40}}\)

OpenStudy (anonymous):

Oh:) lol okay. More?

OpenStudy (anonymous):

Makes band played they did in songs at three current paries. At the first partie they played 1/5 as long as they did at the second partys,which was 1 1/4 as long as the third party. If they played for 5 minutes at the first party, how many minutes did they play at the third party.

OpenStudy (anonymous):

Jakes band played songs at three different partys. That's how it starts outnsorry

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (anonymous):

@RaeDeSol

OpenStudy (anonymous):

@IMStuck

OpenStudy (imstuck):

Your goal here is to solve for h. The first thing they are wanting you to do in order to facilitate this process is to isolate the term with the h in it. You do that by getting rid of the -2/3 that is with the h on the left. You do that by adding it to both sides. What do you get when you do that?

OpenStudy (anonymous):

20 minutes 26 1/4 2831 1/4

OpenStudy (imstuck):

OMG I didn't see that you had several posts in this same question. I was answering the very first one! So sorry!

OpenStudy (anonymous):

28 31 1/4

OpenStudy (anonymous):

Its alright! But for the last ? Those are the options

OpenStudy (anonymous):

28 31 1/4

OpenStudy (imstuck):

Ok, they are telling you that 5 minutes is the length of the playing time for the first party. They also tell you that that 5 minutes is 1/5 of the playing time at the second party. So if 5 minutes is 1/5 of the second party, which is x, our equation for that part is 5=(1/5)x, with x being the length of playing time at the second party. Solve that for x, and you get that they played for 25 minutes at the second party. Now the third party is 1 1/4 shorter than the second party.

OpenStudy (anonymous):

So 26 1/4?

OpenStudy (imstuck):

I'm a bit stuck here...let me think for a minute.

OpenStudy (imstuck):

Ok I got it sorry.

OpenStudy (imstuck):

The third equation is this: The time for the second party is 1 1/4 of the time for the third party. Which translates in numbers and symbols to this: 25=5/4x, with x being the time for the third party. Solve for x and you get 20 minutes.

OpenStudy (anonymous):

Answer options: 20 minutes 26 1/4 28 minutes 31 1/4

OpenStudy (imstuck):

I told you I got it. It's 20 minutes. Ok?

OpenStudy (anonymous):

Yep you got it! Thank you thank yu

OpenStudy (imstuck):

You're welcome!

OpenStudy (anonymous):

Umm could you help with more?

OpenStudy (anonymous):

How could you out eliminate the fractions to make solving then equation 2/16m + 3/16 = 9/16 easier? Divide both sides of the equation by 2/16 Divide both sides if the equation by 16 Multiply both sides of the equation by 16 Subtract 3/16 from both sides of the equation

OpenStudy (imstuck):

That doesn't make any sense, what you wrote. Could you rewrite it, please?

OpenStudy (imstuck):

Sorry, jumped ahead of myself and had to delete that reply...it was incorrect. The correct response is that you subtract 3/16 from both sides, which is the last choice.

OpenStudy (imstuck):

That isolates the term with the unknown in it.

OpenStudy (anonymous):

Ts okay and your a lifesaver

OpenStudy (imstuck):

Thank you for that compliment!

OpenStudy (anonymous):

Welcome!